OpenStudy (anonymous):

what is the soultion set for the equation below? m squared+5m-24=0

OpenStudy (anonymous):

\[m ^{2}+5m-24=0\]

OpenStudy (radar):

m=3, however their is another answer.

OpenStudy (radar):

Use quadratic equation

OpenStudy (anonymous):

i need to find the points so do i plug 3 into all the m's

OpenStudy (anonymous):

also how do you find m

OpenStudy (radar):

Using the quadratic equation: A=1, B=5 and C=-24

OpenStudy (anonymous):

ok hold on let me do that first

OpenStudy (radar):

Hey maybe it will factor (m+8)(m-3)=0 Thats easier than a quadratic solution set is m=3 and m=-8

OpenStudy (radar):

Did you follow that?

OpenStudy (anonymous):

yes let me finish this and i will find another problem

OpenStudy (radar):

I've got to run. On the quadratic equations, first see if it factors if not then go to the quadratic formula.

OpenStudy (anonymous):

ok thank so much

OpenStudy (anonymous):

what quantity should be added to both sides of thi equation tomcomplete the sqares?\[4x ^{2}-10x=3\] a.-25 b.-3 c.25/4 d.25/16

OpenStudy (radar):

d. 25/16

OpenStudy (radar):

When you have been instructed to solve by a certain method such as "completing the squares. Using proper algebraic operations get the equation into this form: Ax^2 + Bx = C Divide thru by A to get X^2 + B/A x=C/A Add (B/A)^2 to both sides; .......will get back with you later, gotta run